Could you help me in finding the perimeter of a right triangle, with only the information of the length of one leg and its area (leg a=18 m, area=63 sq m)?
Answer by richwmiller(17219)

we know the area is 63 sq meters/miles?
one leg is 18
what can we find out if we know the area and one leg?
what is the formula for the area of a triangle?
a=1/2 b*h
the one leg can be either a base or height. We don't care which it is.
63=1/2*18*h
63=9*h
7=h
so now we have the legs of the right triangle.
What can we do with that?
we can find the hypotenuse.
a^2+b^2=c^2
7^2+18^2=c^2
373=c^2
sqrt(373)=c
c=19.3132
so now we have the three sides 18, 7 and 19.3132
add them up to get the perimeter